OpenStudy (anonymous):

An equation is shown below: 2 over 3 multiplied by x plus 8 over 3 equals 2 to the power of x What is the solution to the equation? A. x = –2 B. x = –1 C. x = 1 D. x = 2

OpenStudy (mokeira):

\[\frac{ 2 }{ 3 } \times x+\frac{ 8 }{ 3 }=2^{x}\] is this what you mean? @IndianBoyNiko

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@Mokeira

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x+\frac{ 8 }{ 3 }=2^x\]

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

let me try

OpenStudy (anonymous):

ok

OpenStudy (mokeira):

@ganeshie8

OpenStudy (anonymous):

@Phi

OpenStudy (mokeira):

multiply everything by 3 to remove denominator \[2x+2^{3}=(2^{x})3\]

OpenStudy (mokeira):

that is step one lol

OpenStudy (anonymous):

I got B but I don't know if I'm right :/

OpenStudy (mokeira):

x=2...let me show you how i reasoned although i am not so sure

OpenStudy (anonymous):

okok

OpenStudy (mokeira):

it is same as \[\frac{ 1 }{ 2x }^{-1}+2^{3}=(2^{x})3\] x is found on 2's so we take there exponents as the bases are the same \[-1+3=x\] x=2 you can confirm by substituting

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

You verrrry welcome